OpenStudy (anonymous):

Modify the procedure of Gauss to find the sum. 1+2+3+4+...+155

OpenStudy (tkhunny):

Well, do it... 1+155 = 156 2+154 = 156, and etc.

OpenStudy (anonymous):

$$\ \ S=\ \ \ 1+\ \ \ \ 2+\ \ \ 3+\dots+153+154+155\\\ \ S=155+154+153+\dots+\ \ \ 3+\ \ \ 2+\ \ \ 1\\2S=\underbrace{156+156+156+\dots+156+156+156}_{155 \text{ terms}}\\2S=155(156)\\\ \ S=\frac{155(156)}2\\\ \ S=12090$$

OpenStudy (anonymous):

Thank you oldrin.bataku, that was correct, I finally figured out how to do it the way you have shown above.